Terri wants to make a triangle on her driveway with one side 8ft, one side 17 ft, and one side 8ft. Write an indirect proof to show that Terri’s design is not possible. So far I have that she can't make a RIGHT triangle because of the Pythagorean theorem because a^2 +b^2 = c^2 and 64 + 289 =/= 64, but I'm not sure how to prove it wouldn't work for other types of triangles.
Answer by richwmiller(17219)

Wrong approach
The sum of any two sides of a triangle must be more than the third side.
8+8=16 which is less than 17
BTW The hypotenuse is the largest side of a right triangle
So your right triangle should have been
8^2+8^2=17^2